The demand equation for a certain product is given by
p=100−0.045x, where
p is the unit price (in dollars) of the product and
x is the number of units produced. The total revenue obtained by producing and selling
x units is given by R=xp
.
Determine prices
p
that would yield a revenue of 5910 dollars.
Lowest such price =

Highest such price =

Answer by ankor@dixie-net.com(22740)

The demand equation for a certain product is given by p=100−0.045x, where p is the unit price (in dollars) of the product and x is the number of units produced.
The total revenue obtained by producing and selling
x units is given by R=xp.
Determine prices p that would yield a revenue of 5910 dollars.
xp = 5910
replace p with 100-.045x, from the given equation
x(100-.045x) = 5910
100x - 045x^2 = 5910
arrange as a quadratic equation
-.045x^2 + 100x - 5910 = 0
Using the quadratic equation I got two solutions
x = 2161.46 ~ 2162 units
x = 60.76 ~ 61 units
:
Lowest such price = 100 - .045(2162) = $2.71
:
Highest such price = 100 - .045(62) = $97.21